# Efficient Reasoning Models: A Survey

### v1

@@ 6 Conclusion
In conclusion, this survey provides a comprehensive overview of efficient reasoning techniques. We categorize current efforts into three main directions—shorter, smaller, and faster—each addressing reasoning efficiency from a unique perspective: compressing reasoning chains, building small language models with strong reasoning abilities, and accelerating the decoding stage. We further emphasize the role of model compression in achieving efficient reasoning. As reasoning efficiency continues to gain traction, we believe it holds significant promise for enabling scalable and practical deployment of reasoning models across diverse applications, from real-time systems to resource-constrained environments. We hope this survey serves as a valuable foundation for future research and development in this critical and rapidly evolving field.
@@ Appendix A Appendix
A.1 Overthinking Example
We provide an example to show the overthinking problem: for the simple question “What is the answer of 1 plus 2?”, DeepSeek R1 generates a lengthy response containing 619 tokens, whereas the ideal answer “The answer is 3.” requires only 17 tokens. This highlights the inefficiency brought by overthinking in current models, significantly when solving simple problems.
@@ A.1 Overthinking Example
A.2 Metric Formulas
A.2.1 Carbon Emission
@@
A.2.2 Pass@k
[…]
@@ A.2.2 Pass@k
A.2.4 G-Pass@k
[…]
@@ A.2.4 G-Pass@k
[…]
A.2.5 Outcome and Process Efficiency Metric
@@ A.2.7 Underthinking Metric
A.2.8 Accuracy Efficiency Score
@@
where $\alpha>0$, $\beta>0$, and $\gamma>0$ are weighting factors. Default values $\alpha=1$, $\beta=3$, and $\gamma=5$ are used to emphasize penalizing accuracy drop more heavily than rewarding accuracy improvement.
A.3 Complete List of Datasets and Benchmarks
@@
A.4 Supplementary Results
A.4.1 Results on the AIME dataset
Table 5: Performance of efficient reasoning methods on the AIME dataset. $\dagger$ denotes the result of the original model, averaged over 5 independent runs.
Category Type Methods Acc. / #Tokens Base Model
Original Model - Baseline 70.67% / 10024 DeepSeek-R1-32B
[…]
Shorter KD CoT-Valve 43.30% / 4630 QwQ-32B-Preview
Shorter KD TOPS 46.00% / 6427 Qwen2.5-32B
[…]
Smaller KD DLCoT 53.30% / 18825 Qwen2.5-14B
Smaller RL Open-RS 46.70% / - DeepSeek-R1-Distill-Qwen-1.5B
Smaller RL DeepSacleR 43.10% / - DeepSeek-R1-Distill-Qwen-1.5B
[…]
Faster Efficient sampling $\phi$-Decoding 16.67% / - LLaMA3.1-8B-I
Table 5 presents the performance of various efficient reasoning methods on the AIME dataset. We also include the result of the original model for reference.

### v2

@@ 6 Conclusion
In conclusion, this survey provides a comprehensive overview of efficient reasoning techniques. We categorize current efforts into three main directions—shorter, smaller, and faster—each addressing reasoning efficiency from a unique perspective: compressing reasoning chains, building small language models with strong reasoning abilities, and accelerating the decoding stage. As reasoning efficiency continues to gain traction, we believe it holds significant promise for enabling scalable and practical deployment of reasoning models across diverse applications, from real-time systems to resource-constrained environments. We hope this survey serves as a valuable foundation for future research and development in this critical and rapidly evolving field.
Acknowledgments
This project is supported by the Ministry of Education, Singapore, under its Academic Research Fund Tier 2 (Award Number: MOE-T2EP20122-0006).
@@ Appendix A Appendix
A.1 Details for Model Compression
Quantization.
Quantization improves model efficiency and reduces memory usage by lowering the bit precision of parameters. It is typically categorized into post-training quantization (PTQ) and quantization-aware training (QAT), distinguished by whether retraining is involved. PTQ applies quantization directly to a pre-trained model, while QAT includes a retraining stage to mitigate quantization-induced errors. Quantization can target weights, activations, or both. Advanced methods such as GPTQ (Frantar et al., 2023a), AWQ (Lin et al., 2024), and SmoothQuant (Xiao et al., 2023) further enhance quantization for large language models by reducing activation outliers and minimizing calibration errors.
Pruning.
Pruning reduces model size and inference latency by eliminating redundant or less important parameters. It can be broadly categorized into unstructured pruning, structured pruning, and semi-structured pruning. Unstructured pruning removes individual weights based on certain criteria, such as magnitude. While it achieves high sparsity, it is often less hardware-friendly due to irregular sparsity patterns. Structured pruning eliminates entire units such as neurons, channels, or attention heads, leading to more regular sparsity patterns that are easier to accelerate in practice. Semi-structured pruning strikes a balance between the two, applying constraints such as N:M sparsity, where only a fixed number of weights are retained in each block. This enables efficient execution on specialized hardware. Recent works (e.g., LLM-Pruner, DepGraph) (Ma et al., 2023; Fang et al., 2024; 2023; Feng et al., 2024b), and methods based on importance scores and gradient sensitivity (Wang et al., 2021; Zhang et al., 2021; Tuo & Wang, 2025) have significantly improved the effectiveness and usability of pruning for large models.
Knowledge Distillation.
Knowledge Distillation (KD) transfers the behavior of a large, well-performing teacher model to a smaller student model by aligning output distributions (e.g., logits or soft labels), intermediate representations, or attention patterns. KD approaches can be categorized as black-box or white-box, depending on whether the student has access only to the teacher’s outputs or to internal states as well. Variants like Self-Instruct KD (Wang et al., 2022b) enable the student to inherit reasoning abilities and generalization skills in more flexible settings.
A.2 Overthinking Example
We provide an example to show the overthinking problem: for the simple question “What is the answer of 1 plus 2?”, DeepSeek R1 generates a lengthy response containing 619 tokens, whereas the ideal answer “The answer is 3.” requires only 17 tokens. This highlights the inefficiency brought by overthinking in current models, especially when solving simple problems.
@@ A.2 Overthinking Example
A.3 Latency Comparison of Efficient Reasoning Methods
Table 5 summarizes representative efficient reasoning methods on GSM8K across different categories, providing a practical overview of efficient reasoning approaches for users.
Table 5: Overview of efficient reasoning methods on GSM8K. The speedup ratio is computed mainly through latency comparison, except for Self-Calibration, where sampling count (S.) is used as a proxy.
Category / Type Methods Training Scheme Accuracy Base Model Speedup
Shorter / Routing Self-REF SFT (LoRA) 81.60% LLaMA3-8B-I 1.3 $\times$
Smaller / KD SKIntern Distillation (LoRA) 62.50% LLaMA3-8B-I -
Faster / Efficient self-consistency Path-Consistency Training-free 67.80% LLaMA3-8B-I 1.2 $\times$
Shorter / SFT CoT-Valve Progressive SFT (LoRA) 87.30% LLaMA3.1-8B-I 1.7 $\times$
Shorter / SFT TokenSkip SFT (LoRA) 78.20% LLaMA3.1-8B-I 1.7 - 1.8 $\times$
Shorter / SFT TALE-PT SFT (LoRA) 78.57% LLaMA3.1-8B-I 1.7 $\times$
Shorter / Latent reasoning SoftCoT SFT (Freeze FT) 81.03% LLaMA3.1-8B-I 4.0 - 5.0 $\times$
Shorter / Latent reasoning LightThinker SFT (Full FT) 88.25% LLaMA3.1-8B-I up to 1.4 $\times$
Shorter / Latent reasoning Token Assorted SFT (Full FT) 84.10% LLaMA3.1-8B-I 1.2 $\times$
Smaller / KD Mix Mixed distillation (Full FT & LoRA) 81.40% LLaMA3.1-8B-I -
Smaller / KD DLCoT Distillation (Full FT) 93.60% LLaMA3.1-8B-I -
Faster / Efficient sampling $\phi$-Decoding Training-free 86.58% LLaMA3.1-8B-I 2.8 $\times$
Faster / Efficient self-consistency Self-Calibration SFT (Full FT) 80.43% LLaMA3.1-8B-I 16.7 $\times$ (S.)
A.4 Metric Formulas
A.4.1 Carbon Emission
@@
A.4.2 Pass@k
$\text{Pass@}k=1-\mathbb{E}_{\text{task}}\left[\frac{{n-c\choose k}}{{n\choose k}}\right$] (2)
@@ A.4.2 Pass@k
A.4.4 G-Pass@k
$\text{G-Pass@}k_{\tau}=\mathbb{E}_{\text{task}}\left[\sum_{j=\lceil\tau k\rceil}^{c}\frac{{c\choose j}{n-c\choose k-j}}{{n\choose k}}\right$] (4)
@@ A.4.4 G-Pass@k
$\text{mG-Pass@}k_{\tau}=\frac{2}{k}\sum_{i=\lceil 0.5k\rceil+1}^{k}\text{G-Pass@}k_{\frac{i}{k}}$ (5)
A.4.5 Outcome and Process Efficiency Metric
@@ A.4.7 Underthinking Metric
A.4.8 Accuracy Efficiency Score
@@
where $\alpha>0$, $\beta>0$, and $\gamma>0$ are weighting factors. The default values $\alpha=1$, $\beta=3$, and $\gamma=5$ are used to emphasize penalizing accuracy drop more heavily than rewarding accuracy improvement.
A.5 Complete List of Datasets and Benchmarks